OpenStudy (anonymous):

can someone help me please Write an equation of a line in point slope form that has a slope of -3 and passes through the point (3, -4).

OpenStudy (anonymous):

y-y1=m(x-x1) y-(-4)=-3(x-3) y+4=-3x+9 y=-3x+5

OpenStudy (anonymous):

can u help me with more if i post in new discussions ? @03225186213

OpenStudy (anonymous):

yes why not? but you should it first then i will chek it?